Two identical machines are positioned the same distance from a worker. The intensity of sound delivered by each machine at the location of the worker is 2.0 × 107 W/m2.
Find the sound level heard by the worker:
a) when one machine is operating
b) when both machines are operating.
1
Expert's answer
2020-04-17T10:51:31-0400

a)


"\beta_1=10\lg{\frac{2\cdot 10^{-7}}{10^{-12}}}=10\lg{(2\cdot 10^{5})}=53\ dB"

b)


"\beta_2=10\lg{\frac{4\cdot 10^{-7}}{10^{-12}}}=10\lg{(4\cdot 10^{5})}=56\ dB"